A circular copper waveguide has a radius of 0.4 cm and is filled with a dielectric material having 𝜀𝑟 = 1.5 and tan δ = 0.0002.

2.1.1. Identify the first four propagating modes and their cut-off frequencies.

2.1.2. For the dominant mode, calculate the attenuation due to dielectric loss and attenuation due to conductor loss at 20 GHz.


1
Expert's answer
2022-03-03T12:05:06-0500

1)

"\nu_1=20~GHz,"

"\nu_2=5.0~GHz,"

"\nu_3=2.2~GHz,"

"\nu_4=1.3~GHz."

2)

"k_d=\omega\sqrt{\mu\epsilon(1-j\tan\delta)}=0.4~GHz,"

"k_c=\beta-j\alpha=350~MHz."